# Synchronization-induced violation of thermodynamic uncertainty relations

### v1

@@ II.4 TUR violation for local power
where $Q_{P}=\Omega\coth(\Omega/(2T))/T$. Since $Q_{P}\geq 2$ (see Appendix C), violations of the TUR for $Q_{P_{A}}$, if any, must originate from the last fraction in the r.h.s. of the equation. In particular, one is interested in large values of the denominator, and hence one should look for $\delta P_{A}/P_{A}^{(0)}\gg 1$.
In the adiabatic regime ($\Omega\ll\omega_{l},T$) one has
$P_{l}^{(0)}\to\Omega^{2}\alpha_{l}^{({\rm ad})},$
$\delta P_{l}\to\Omega\delta\alpha^{({\rm ad})}_{l}~{},$ (20)
@@
and hence $P\propto\Omega^{2}$ and $P_{l}\propto\Omega$ at leading order. Here, we have introduced auxiliary functions $\alpha_{l}^{({\rm ad})}$ and $\delta\alpha_{l}^{({\rm ad})}$ that do not depend on the external frequency $\Omega$ anymore (see Appendix D). This leads to the adiabatic expansion for the local TUR:
$Q^{({\rm ad})}_{P_{A}}=2\Omega^{2}(\alpha_{A}^{({\rm ad})}+\alpha_{B}^{({\rm ad})})\frac{\alpha_{A}^{({\rm ad})}}{{\delta\alpha_{A}^{({\rm ad})}}^{2}}~{}.$ (21)
From the above expression, one expects a violation of the local TUR ($Q_{P_{A}}\leq 2$) at sufficiently small $\Omega$, regardless of the precise shape of the spectral density ${\cal J}_{1}(\omega)$. This is indeed the case, as shown in Fig. 3(a-b).
@@
Figure 3: Quantifying fluctuations of $P_{A}$ via the trade-off quantity $Q_{P_{A}}$ in the isothermal regime. Density plots of $Q_{P_{A}}$ (a) and associated power $P_{A}$ (c) as a function of $\Omega/\omega_{A}$ and $T/\omega_{A}$ in the large cut-off regime $\omega_{c}=1000\omega_{A}$. Panels (b,d) report the same quantities as in panels (a,c), respectively, but for a small cut-off $\omega_{c}=1.2\omega_{A}$. All results have been obtained in the strong damping regime $\gamma_{2}=100\omega_{A}$; other parameters values are $\omega_{B}=0.6\omega_{A}$, $\gamma_{1}=0.01\omega_{A}$, and $\phi=\pi/2$. In panels (a,b) the dashed line denotes $Q_{P_{A}}=2$ separating regions where the TUR is violated (cool colors, $Q_{P_{A}}<2$) from regions where it is not (warm colors, $Q_{P_{A}}>2$). In panels (c,d) the dashed line refers to $P_{A}=0$ separating negative power contributions (power production, cool colors) from positive ones (power absorption, warm colors).
@@
There, density plots of $Q_{P_{A}}$ in the $\Omega-T$ plane are reported for two different values of $\omega_{c}$, where a region with $Q_{P_{A}}<2$ is clearly visible in the left corner corresponding to $\Omega<\omega_{A}$. Looking at the corresponding power contribution $P_{A}$ (see panels c-d), one finds that $P_{A}<0$ in almost the same region, i.e., the system is acting as a work-to-work converter. This demonstrates that in the strong damping regime it is possible to obtain deviations from the TUR bound in the subsystem performance due to the breaking of TRS. However, violation of the TUR in this parameter regime is linked to small power magnitude, with $P_{A}\to\Omega{\delta\alpha}^{({\rm ad})}_{A}$ in the adiabatic limit. In passing we mention that for the $l=B$ channel at $\phi=\pi/2$ analogous TUR violation are observed but always with $P_{B}>0$ values, that is $l=B$ cannot be used as a useful resource for power production (see Supplementary Note 1). One may thus wonder if it is possible to achieve precise ($Q_{P_{A}}\ll 2$) but sizeable local power signals. To this end, the opposite, diabatic, regime ($\Omega\gg\omega_{l},\omega_{c}$) can be inspected. There, one obtains the following asymptotic expansions (see Appendix D)
$P_{l}^{(0)}\to\Omega{\cal J}_{1}(\Omega)\alpha_{l}^{({\rm dia})},$
$\delta P_{l}\to\frac{\Omega^{2}}{\omega_{c}}{\cal J}_{1}(\Omega)\delta\alpha_{l}^{({\rm dia})}~{},$ (22)
with the total power $P\propto\Omega{\cal J}_{1}(\Omega)$. Clearly, these expressions, and hence the behaviour of Eq. (19), depend on the shape of ${\cal J}_{1}(\omega)$ and on the value of the cut-off $\omega_{c}$. Indeed, one gets
$Q_{P_{A}}^{({\rm dia})}=\frac{\Omega}{T}\frac{1+\alpha_{A}^{({\rm dia})}/\alpha_{B}^{({\rm dia})}}{\left[1+\tilde{{\cal C}}\frac{\Omega}{\omega_{c}}\right]^{2}},$ (23)
@@
where $\tilde{{\cal C}}=\delta\alpha_{A}^{({\rm dia})}/\alpha_{A}^{({\rm dia})}$. First of all, if one considers $\omega_{c}$ as the largest energy scale, thus with no memory effects, no TUR violations are expected (see the rightmost regions in the density plot of Fig. 3(a)). Intriguingly, the situation is different in the case of small cut-off $\omega_{c}$, i.e., when non-Markovian effects become important. Indeed, in the case of small $\omega_{c}$ and considering finite values of $\tilde{{\cal C}}$ one thus gets $Q_{P_{A}}^{({\rm dia})}\propto 1/\Omega$, which again shows the possibility to get $Q_{P_{A}}<2$. To corroborate this finding, in Fig. 3(b) the density plot of $Q_{P_{A}}$ in the $\Omega-T$ plane is reported for a representative small value $\omega_{c}=1.2\omega_{A}$. In this figure two regions where $Q_{P_{A}}<2$ are present: the first in the adiabatic regime, as discussed above, and a second new region in the diabatic regime $\Omega\gg\omega_{l},\omega_{c}$. The appearance of this new region is tightly bound to memory effects due to a finite cut-off $\omega_{c}$. More importantly, this latter regime is linked to finite power magnitude. This is indeed shown in Fig. 3(d), where a region with negative power contribution $P_{A}<0$ with sizeable magnitude is evident. Physically, in the diabatic regime $\Omega\gg\omega_{l},\omega_{c}$, the external frequency is much higher than the cut-off frequency of the reservoir that becomes effectively freezed, and the large amount of injected power from the $l=B$ channel is almost entirely transferred to the $l=A$ one, resulting in a very efficient work-to-work conversion with efficiency $\eta\sim 1$ (see Supplementary Note 2). We stress that in our model, in addition to TRS breaking, two are the key ingredients to achieve local TUR violation together with finite output power: memory effects and synchronization induced by strong damping $\gamma_{2}$. Interestingly, in the diabatic regime $\Omega\gg\omega_{l},\omega_{c}$, a quantity reminiscent of the Pearson coefficient of Eq. (10) naturally appears. Indeed, looking at the denominator in Eq. (23) one has $1+\tilde{{\cal C}}\Omega/\omega_{c}$. Inspecting the definition of $\tilde{{\cal C}}$ one finds that
@@
and it is $|\tilde{{\cal C}}|<1$. The behaviour of this Pearson-like coefficient is qualitatively the same as the one reported in Fig. 2(b) for the Pearson coefficient ${\cal C}$. At low temperature $\tilde{{\cal C}}$ reaches values close to $-1$ in the strong damping regime when the two subsystems reach full synchronization being in anti-phase. This allows to get large values of the denominator in Eq. (23) that is a necessary condition to achieve TUR violation with sizeable power, clearly showing the importance of synchronization, established at strong damping. In Supplementary Note 3 for the sake of completeness we have reported the behaviour of the TUR and the subsystem power in the case of weak damping $\gamma_{2}$, where synchronization is lacking, showing that there the TUR for ${P_{A}}$ is not violated in the diabatic regime.
@@
Figure 4: Quantifying fluctuations of $P_{A}$ via the trade-off quantity $Q_{P_{A}}$, in the presence of a temperature gradient in the strong damping regime $\gamma_{2}=100\omega_{A}$, with cut-off $\omega_{c}=1.2\omega_{A}$. Density plots of (a) $Q_{P_{A}}$ and (b) $P_{A}$ as a function of the driving frequency $\Omega/\omega_{A}$ and average temperature $\bar{T}/\omega_{A}$ at relative temperature gradient $\Delta T/\bar{T}=1.8$. Other parameters as in Fig. 3.
@@
So far we have discussed results in the isothermal regime with $T_{1}=T_{2}=T$. Before closing, we now demonstrate the robustness of our results in the presence of a relative temperature gradient $\Delta T/\bar{T}$, with $\bar{T}=(T_{1}+T_{2})/2$ the average temperature and $\Delta T=T_{1}-T_{2}$ the temperature difference. Not only the presence of a region of TUR violation with sizeable power in the diabatic regime still holds with a finite temperature gradient, but also wider regions are obtained in the non-linear regime of temperature gradient. In Fig. 4(a,b) we show results for $Q_{P_{A}}$ and $P_{A}$, respectively, as a function of $\Omega$ and $\bar{T}$, at given $\Delta T/\bar{T}=1.8$. Notice that this last value implies a strong unbalance $T_{2}\ll T_{1}$ (strongly non-linear regime). At low driving frequencies, $\Omega<\omega_{A}$, the TUR for $P_{A}$ is violated for very low average temperature, $\bar{T}\lesssim 0.1\omega_{A}$. Instead, at higher driving frequencies the TUR is violated almost independently of the average temperatures considered, $0<\bar{T}<2\omega_{A}$. Comparing Fig. 4(b) and Fig. 3(b), we can see that the temperature gradient between the hot bath $\nu=1$ and the cold one $\nu=2$ can be exploited to greatly lower the driving frequency required to violate the TUR for $P_{A}$ (see also Supplementary Note 4).
III Discussion
@@
Our results show the intimate connection between synchronization and thermodynamics of precision. From the standpoint of quantum technologies, synchronization mechanisms could be exploited to obtain finite and precise power in quantum circuits, where non-Markovian environments can be engineered [85, 84, 86, 87]. From a more general point of view, a thermodynamic perspective could also be useful in the broad field of classical synchronization.

### v2

@@ IV TUR violation for local power
where $Q_{P}=\Omega\coth\left[\Omega/(2T)\right]/T$. Since $Q_{P}\geq 2$ (see Appendix C), violations of the TUR for $Q_{P_{A}}$, if any, must originate from the last fraction in the r.h.s. of the equation. In particular, one is interested in large values of the denominator, and hence one should look for $\delta P_{A}/P_{A}^{(0)}\gg 1$.
It is thus instructive to study the two opposite asymptotic behaviours of small and large external frequency. In the adiabatic regime ($\Omega\ll\omega_{l},T$) one has
$P_{l}^{(0)}\to\Omega^{2}f_{l}^{2}(\Omega)\alpha_{l}^{({\rm ad})},$
$\delta P_{l}\to\Omega f_{B}(\Omega)\delta\alpha^{({\rm ad})}_{l}~{},$ (21)
@@
where we have indicated with $\alpha_{l}^{({\rm ad})}$ and $\delta\alpha_{l}^{({\rm ad})}$ the expansion coefficients that do not depend on the external frequency $\Omega$ anymore (see Appendix D for their explicit expressions). This leads to the adiabatic expansion for the local TUR quantifier:
$Q^{({\rm ad})}_{P_{A}}=2\Omega^{2}\left(\alpha_{A}^{({\rm ad})}+f_{B}^{2}(\Omega)\alpha_{B}^{({\rm ad})}\right)\frac{\alpha_{A}^{({\rm ad})}}{f_{B}^{2}(\Omega){\delta\alpha_{A}^{({\rm ad})}}^{2}}~{}.$ (22)
In the opposite, diabatic, regime ($\Omega\gg\omega_{l},\omega_{c}$) one obtains the following asymptotic expansions
$P_{l}^{(0)}\to\Omega{\cal J}_{1}(\Omega)f^{2}_{l}(\Omega)\alpha_{l}^{({\rm dia})},$
$\delta P_{l}\to\frac{\Omega^{2}}{\omega_{c}}f_{B}(\Omega){\cal J}_{1}(\Omega)\delta\alpha_{l}^{({\rm dia})}~{},$ (23)
where again $\alpha_{l}^{({\rm dia})}$ and $\delta\alpha_{l}^{({\rm dia})}$ are frequency-independent expansion coefficients (see Appendix D). These expressions depend on the shape of ${\cal J}_{1}(\Omega)$ and on the value of the cut-off $\omega_{c}$. The corresponding TUR quantifier becomes
$Q_{P_{A}}^{({\rm dia})}=\frac{\Omega}{T}\frac{1+f_{B}^{2}(\Omega)\alpha_{B}^{({\rm dia})}/\alpha_{A}^{({\rm dia})}}{\left[1+\tilde{{\cal C}}f_{B}(\Omega)\frac{\Omega}{\omega_{c}}\right]^{2}},$ (24)
where $\tilde{{\cal C}}=\delta\alpha_{A}^{({\rm dia})}/\alpha_{A}^{({\rm dia})}$. Due to the above scaling behaviours, in both the asymptotic regimes one gets the possibility to achieve TUR violations as we will now discuss.
IV.1 Standard TUR violation
Let us now start by discussing possible violations from the standard TUR $Q_{P_{A}}\geq 2$. To this end, we first consider the simple case of $r=0$ that corresponds to equal weigths $f_{A}(\Omega)=f_{B}(\Omega)=1$. In the adiabatic regime, from Eq. (22) one gets a $Q_{P_{A}}^{({\rm ad})}\propto\Omega^{2}$ regardless of the precise shape of the spectral density ${\cal J}_{1}(\omega)$. This is indeed the case, as shown in figure 3(a-b)
@@ IV.1 Standard TUR violation
Figure 3: Quantifying fluctuations of $P_{A}$ via the trade-off quantity $Q_{P_{A}}$ in the isothermal regime. Density plots of $Q_{P_{A}}$ (a) and associated power $P_{A}$ (c) as a function of $\Omega/\omega_{A}$ and $T/\omega_{A}$ for $r=0$ in the large cut-off regime $\omega_{c}=1000\omega_{A}$. Panels (b,d) report the same quantities as in panels (a,c), respectively, but for a small cut-off $\omega_{c}=1.2\omega_{A}$. All results have been obtained in the strong damping regime $\gamma_{2}=100\omega_{A}$; other parameters values are $\omega_{B}=0.6\omega_{A}$, $\gamma_{1}=0.01\omega_{A}$, and $\phi=\pi/2$. In panels (a,b) the dashed line denotes $Q_{P_{A}}=2$ separating regions where the TUR is violated (cool colors, $Q_{P_{A}}<2$) from regions where it is not (warm colors, $Q_{P_{A}}>2$). In panels (c,d) the dashed line refers to $P_{A}=0$ separating negative power contributions (power production, cool colors) from positive ones (power absorption, warm colors).
@@
There, density plots of $Q_{P_{A}}$ in the $\Omega-T$ plane are reported for two different values of $\omega_{c}$, where a region with $Q_{P_{A}}<2$ is clearly visible in the left corner corresponding to $\Omega<\omega_{A}$. Looking at the corresponding power contribution $P_{A}$ (see panels c-d), one finds that $P_{A}<0$ in almost the same region, i.e., the system is acting as a work-to-work converter. However, violation of the standard TUR in this parameter regime is associated with small power magnitude, with $P_{A}\to\Omega{\delta\alpha}^{({\rm ad})}_{A}$ in the adiabatic limit. In passing we mention that for the $l=B$ channel at $\phi=\pi/2$ analogous TUR violation are observed but always with $P_{B}>0$ values, that is $l=B$ cannot be used as a useful resource for power production (see Appendix E).
@@
One may thus wonder if it is possible to achieve precise ($Q_{P_{A}}\ll 2$) but sizeable local power signals. To this end, one can look at the large frequency regime. First of all, if one considers $\omega_{c}$ as the largest energy scale, thus with no memory effects, no TUR violations are expected (see the rightmost regions in the density plot of figure 3(a)). Intriguingly, the situation is different in the case of small cut-off $\omega_{c}$, i.e., when non-Markovian effects become important. Indeed, in the case of small $\omega_{c}$ and considering finite values of $\tilde{{\cal C}}$ one thus gets $Q_{P_{A}}^{({\rm dia})}\propto 1/\Omega$, which again shows the possibility to get $Q_{P_{A}}<2$. To corroborate this finding, in figure 3(b) the density plot of $Q_{P_{A}}$ in the $\Omega-T$ plane is reported for a representative small value $\omega_{c}=1.2\omega_{A}$. In this figure two regions where $Q_{P_{A}}<2$ are present: the first in the adiabatic regime, as discussed above, and a second new region in the diabatic regime $\Omega\gg\omega_{l},\omega_{c}$. Importantly, this latter regime is also associated with finite power magnitude. This is indeed shown in figure 3(d), where a region with negative power contribution $P_{A}<0$ with sizeable magnitude is evident. Physically, in the diabatic regime $\Omega\gg\omega_{l},\omega_{c}$, the external frequency is much higher than the cut-off frequency of the reservoir that becomes effectively freezed, and the large amount of injected power from the $l=B$ channel is almost entirely transferred to the $l=A$ one, resulting in a very efficient work-to-work conversion with efficiency $\eta\sim 1$ (see Appendix F).
We stress that in our model, in addition to TRS breaking, two are the key ingredients to achieve local TUR violation together with finite output power: memory effects and synchronization induced by strong damping $\gamma_{2}$. Indeed, related to the latter point, in Fig. 4 one can notice a connection between the TUR quantifier and the Pearson coefficient ${\cal C}$. In particular, the stronger the synchronization, the lower the value of the TUR quantifier.
Figure 4: Standard TUR violation and synchronization. Parametric plot of the ratio $Q_{P_{A}}/2$ as a function of the Pearson coefficient at varying damping strengths $\gamma_{2}$. The stronger the damping, the more pronounced the (anti)synchronization ($\mathcal{C}\to-1$), the more the TUR is violated ($Q_{P_{A}}/2<1$). Results for $r=0$, $T=0.1\omega_{A}$, $\omega_{c}=1.2\omega_{A}$, and $\Omega=200\omega_{A}$ (diabatic regime). Other parameters are $\omega_{B}=0.6\omega_{A}$, $\gamma_{1}=0.01\omega_{A}$, and $\phi=\pi/2$.
Interestingly, in the diabatic regime $\Omega\gg\omega_{l},\omega_{c}$, a quantity reminiscent of the Pearson coefficient of equation (11) naturally appears. Indeed, looking at the denominator in equation (24) and inspecting the definition of $\tilde{{\cal C}}$ one finds that
@@
and it is $|\tilde{{\cal C}}|<1$. The behaviour of this Pearson-like coefficient is qualitatively the same as the one reported in figure 2(b) for the Pearson coefficient ${\cal C}$. At low temperature $\tilde{{\cal C}}$ reaches values close to $-1$ in the strong damping regime when the two subsystems reach full synchronization being in anti-phase. This allows to get large values of the denominator in equation (24) that is a necessary condition to achieve TUR violation with sizeable power, clearly showing the importance of synchronization, established at strong damping. In Appendix G for the sake of completeness we have reported the behaviour of the TUR and the subsystem power in the case of weak damping $\gamma_{2}$, where synchronization is lacking, showing that there the standard TUR for ${P_{A}}$ is not violated in the diabatic regime.
@@
Figure 5: Quantifying fluctuations of $P_{A}$ via the trade-off quantity $Q_{P_{A}}$, in the presence of a temperature gradient in the strong damping regime $\gamma_{2}=100\omega_{A}$, with cut-off $\omega_{c}=1.2\omega_{A}$ and for $r=0$. Density plots of (a) $Q_{P_{A}}$ and (b) $P_{A}$ as a function of the driving frequency $\Omega/\omega_{A}$ and average temperature $\bar{T}/\omega_{A}$ at relative temperature gradient $\Delta T/\bar{T}=1.8$. Other parameters as in figure 3.
@@
So far we have discussed results in the isothermal regime with $T_{1}=T_{2}=T$. Before closing, we now demonstrate the robustness of our results in the presence of a relative temperature gradient $\Delta T/\bar{T}$, with $\bar{T}=(T_{1}+T_{2})/2$ the average temperature and $\Delta T=T_{1}-T_{2}$ the temperature difference. Not only the presence of a region of standard TUR violation with sizeable power in the diabatic regime still holds with a finite temperature gradient, but also wider regions are obtained in the non-linear regime of temperature gradient. In figure 5(a,b) we show results for $Q_{P_{A}}$ and $P_{A}$, respectively, as a function of $\Omega$ and $\bar{T}$, at given $\Delta T/\bar{T}=1.8$. Notice that this last value implies a strong unbalance $T_{2}\ll T_{1}$ (strongly non-linear regime). At low driving frequencies, $\Omega<\omega_{A}$, the TUR for $P_{A}$ is violated for very low average temperature, $\bar{T}\lesssim 0.1\omega_{A}$. Instead, at higher driving frequencies the TUR is violated almost independently of the average temperatures considered, $0<\bar{T}<2\omega_{A}$. Comparing figure 5(b) and figure 3(b), we can see that the temperature gradient between the hot bath $\nu=1$ and the cold one $\nu=2$ can be exploited to greatly lower the driving frequency required to violate the TUR for $P_{A}$ (see also Appendix H). In closing this Section we mention that considering $r>0$ the standard TUR is still violated and no qualitative changes are found.
IV.2 Generalized TUR for periodic drives
So far we have discussed possible violations from the standard TUR bound $Q_{P_{A}}\geq 2$. However, the authors of Refs. [61, 62, 63], considering a system under time-dependent drives, have derived a generalized bound $Q_{O}\geq{\cal V}_{O}$ with the quantity in the r.h.s. that in our case for the local power $P_{l}$ reads
${\cal V}_{P_{l}}=2\left(1-\Omega\frac{\partial_{\Omega}P_{l}}{P_{l}}\right)^{2}.$ (26)
Focusing again on the $l=A$ channel, and looking for precise but sizeable output power, we now compare the TUR quantifier $Q_{P_{A}}$ to this generalized bound in the diabatic regime ($\Omega\gg\omega_{l},\omega_{c}$). First of all, inspecting Eq. (IV), the asymptotic behaviour of Eq. (26) in the case $r=0$ reads ${\cal V}_{P_{A}}\propto 1/\Omega^{2}$. Recalling that the TUR quantifier $Q_{P_{A}}\propto 1/\Omega$ in the diabatic regime, regardless of the value of the parameter $r$, the generalized TUR bound turns out to be satisfied with $Q_{P_{A}}\geq{\cal V}_{P_{A}}$. However, one can consider a more general situation with $r>0$. Looking at Eq. (IV) this corresponds to optimize the magnitude of the local output power. In this case, the asymptotic behaviour of Eq. (26) becomes ${\cal V}_{P_{A}}\to 2r^{2}$, independent of the external frequency $\Omega$. This allows to look for the simultaneous desiderata of precise ($Q_{P_{A}}/{\cal V}_{P_{A}}\ll 1$), sizeable local power $P_{A}<0$, and high work-to-work conversion efficiency $\eta$ possibly close to unity in the isothermal regime. This last requirement imposes a constraint (see Appendix F) on the value of the parameter, leading to $10<r<$ 1. Indeed, for $r>1$ the work to work efficiency in the diabatic regime tends to vanish. Therefore, by choosing $10<r<$ 1 it is possible to obtain TUR violations, together with finite output power, even considering the generalized TUR bound posed by Eq. (26). In Fig. 6 we explicitly show a density plot of the TUR ratio $Q_{P_{A}}/{\cal V}_{P_{A}}$ for a representative value $r=1/2$, demonstrating the generalized TUR violation, again linked to the regime of synchronization achieved at strong damping, as discussed above.
Figure 6: Violation of the generalized TUR bound. (a) Density plot of $Q_{P_{A}}/{\cal V}_{P_{A}}$ with $r=1/2$ as a function of $\Omega/\omega_{A}$ and $T/\omega_{A}$ for small cut-off $\omega_{c}=1.2\omega_{A}$. Other parameters as in figure 3. (b) Density plot of the local power $P_{A}$ for the same parameters as in Panel (a).
V Discussion
@@
Our results show the intimate connection between synchronization and thermodynamics of precision. From the standpoint of quantum technologies, synchronization mechanisms could be exploited to obtain finite and precise power in quantum circuits, where non-Markovian environments can be engineered [89, 88, 90, 91]. From a more general point of view, a thermodynamic perspective could also be useful in the broad field of classical synchronization. These results open interesting perspectives for the exploitation of synchronization mechanisms, including bath-induced ones [93], beyond the paradigmatic model studied here. For instance it would be interesting to consider the impact of non-linearities and other systems including two-level systems that are commonly used in superconducting circuits.
Acknowledgements.
M.C. and M.S. acknowledge support from the project PRIN 2022 - 2022PH852L (PE3) TopoFlags - ”Non reciprocal supercurrent and topological transition in hybrid Nb-InSb nanoflags” funded within the programme ”PNRR Missione 4 - Componente 2 - Investimento 1.1 Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN)”, funded by the European Union - Next Generation EU. L.R. and G.B. acknowledge financial support from the Julian Schwinger Foundation (Grant JSF-21-04-0001) and from INFN through the project “QUANTUM”. L.R., F.C., and G.B. acknowledge support from the project PRIN 2022 - 2022XK5CPX (PE3) SoS-QuBa - ”Solid State Quantum Batteries: Characterization and Optimization” funded within the programme ”PNRR Missione 4 - Componente 2 - Investimento 1.1 Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN)”, funded by the European Union - Next Generation EU.